OpenStudy (anonymous):

Solve the following system using the substitution method. x + 6y = –5 2x – 9y = –17

hero (hero):

Solve each equation for y: y = (-5-x)/6 y = (17 + 2x)/9 Substitute y = y to eliminate y : (-5-x)/6 = (17 + 2x)/9 Cross Multiply: 9(-5-x) = 6(17+2x) -45 -9x = 102+12x Continue Solving for x: 21x = -147 x = 7 y = -2